# Parameter-Dependent Control Lyapunov Functions for Stabilizing Nonlinear Parameter-Varying Systems

### v1

@@ Abstract
This paper introduces the concept of parameter-dependent (PD) control Lyapunov functions (CLFs) for gain-scheduled stabilization of nonlinear parameter-varying (NPV) systems under input constraints. It shows that given a PD-CLF, a locally Lipschitz control law can be constructed by solving a robust quadratic program. For polynomial control-affine NPV systems, it provides convex conditions, based on the sum of squares (SOS) programming, to jointly synthesize both a PD-CLF and a PD controller, aimed at maximizing the PD region of stabilization. Input constraints can be straightforwardly incorporated into the synthesis procedure. Unlike traditional linear parameter-varying (LPV) methods that rely on linearization or over-approximation to get an LPV model, the proposed framework fully captures the nonlinearities of the system dynamics. Simulation results validate the efficacy of the method, showcasing its potential for stabilizing NPV systems under input constraints.
@@ I Introduction
Statement of Contributions: This paper introduces the concepts of parameter-dependent CLF (PD-CLF) and PD region of stabilization for NPV systems. It shows that given a PD-CLF, a locally Lipschitz control law can be constructed by solving a robust quadratic program. Additionally, it presents a method for jointly synthesizing a PD-CLF and a PD controller while maximizing the PD region of stabilization based on SOS programming. It is also shown that input constraints can be straightforwardly incorporated into the synthesis procedure. Finally, we validate the theoretical development using simulation results.
@@ II Preliminaries
Unlike existing NPV work [16, 17, 18, 19], we do not require $\dot{\theta}$ to be online measurable.
@@ III Parameter-Dependent CLFs for Stabilizing NPV Systems
III-A Constructing control law via robust quadratic programs
@@
$\alpha_{1}(\left\lVert x\right\rVert)\leq V\leq\alpha_{2}(\left\lVert x\right\rVert),\quad\forall t\geq 0,\forall x.$ (14)
@@
IV Local Analysis Under Input Constraints
In this section, we consider local stabilizability established by a PD-CLF. To make the problem more practically meaningful, we also consider input constraints: $u(t)\in\mathcal{U}\subseteq\mathbb{R}^{m}$, where $\mathcal{U}$ denotes the set of admissible inputs and is assumed to be a polytopic set.
@@ IV Local Analysis Under Input Constraints
[…]
$\hskip 8.53581pt\underbrace{\frac{\partial V}{\partial x}f+\frac{\partial V}{\partial x}gu+\frac{\partial V}{\partial\theta}\dot{\theta}}_{\triangleq\dot{V}(x,\theta,\dot{\theta},u)}\leq-\alpha(\left\lVert x\right\rVert),\vspace{-8mm}$ (15b)
@@
A PD region of stabilization (PD-ROS) for system 1 is the set of initial states $x(0)$ and parameters $\theta(0)$ from which there exists a control law $K(x,\theta)$ such that $u(t)=K(x(t),\theta(t))\in\mathcal{U},\ \forall t\geq 0$, and $\lim_{t\rightarrow 0}x(t)=0$ for all admissible trajectories of $\theta$. Formally, a PD-ROS is a set defined by $\left\{(x(0),\theta(0))\in\mathbb{R}^{n}\times\Theta\!:\ \forall\theta\in\mathcal{F}_{\Theta}^{\mathcal{V}},\right.\\ \exists u(t)=K(x(t),\theta(t))\in\mathcal{U}\ \forall t\geq 0,\left.\textup{s.t.}\lim_{t\rightarrow 0}x(t)=0.\right\}$
@@
A robust region of stabilization (R-ROS) for system 1 is the set of initial states $x(0)$ from which there exists a control law $K(x,\theta)$ such that $u(t)=K(x(t),\theta(t))\in\mathcal{U}$ $\forall t\geq 0$, and $\lim_{t\rightarrow 0}x(t)=0$ for all admissible trajectories of $\theta$. Formally, an R-ROS is a set defined by $\left\{x(0)\in\!\mathbb{R}^{n}\!\!:\ \forall\theta\in\mathcal{F}_{\Theta}^{\mathcal{V}},\ \exists u(t)=K(x(t),\theta(t))\in\mathcal{U}\ \forall t\geq 0,\right.\\ \left.\textup{s.t.}\lim_{t\rightarrow 0}x(t)=0.\right\}$
@@
Compared to R-ROS, introducing the concept of PD-ROS allows us to enlarge the set of states from which the system can be stabilized by explicitly considering $\theta(0)$, the initial value of $\theta$. This is demonstrated by Fig. 2 to be presented in Section VI.
@@
Condition 15b ensures that if $(x(0),\theta(0))\in\Omega(V,\rho)$, then there exist $u(t)\in\mathcal{U}$ $\forall t\geq 0$ such that
@@
V Convex synthesis of PD-CLFs Using SOS Programming
@@
Note that a polynomial system 1 can always be written in the form of 1 if $f(x,\theta)$ does not contain terms independent of $x$, i.e., monomials of zero degrees in $x$. When $f(x,\theta)$ contains $x$-independent terms, state transformation may be leveraged to obtain a linear-like form, as we will demonstrate in Section VI-B.
@@
V-A Preliminaries about SOS optimization
@@
(S-procedure) [23] Let $p(x)$ be a a polynomial and let $\mathcal{X}$ be a set defined in 20. Suppose there exist SOS polynomials $\lambda_{i}(x)$, $i=0,1,\dots,n$ such that $\left(1+\lambda_{0}(x)\right)p(x)-\sum_{i=1}^{n}\lambda_{i}(x)c_{i}(x)\in{\Phi_{\textup{sos}}(x)}$. Then, $p(x)\geq 0\textup{ for all }x\in\mathcal{X}$.
@@
[22, Proposition 10] Let $F(x)$ be an $N\times n$ symmetric polynomial matrix of degree $2d$ in $x\in\mathbb{R}^{n}$, $\mathcal{X}$ is a semialgebraic set defined in 20, and $v\in\mathbb{R}^{N}$. Suppose there exist SOS polynomials $\lambda_{i}(x,v)$, $i=1,\dots,n$ such that $v^{T}F(x)v-\sum_{i=1}^{n}\lambda_{i}(x,v)c_{i}(x)\in{\Phi_{\textup{sos}}(x,v)}$. Then, $F(x)\geq 0\textup{ for all }x\in\mathcal{X}$.
V-B Convex Synthesis of a PD-CLF and a PD Controller
@@
Consider an NPV system 18 with the parameters $\theta$ satisfying 2, and the set $\mathcal{X}$ defined in 20. Suppose there exists polynomial matrices $X(\tilde{x},\theta)\in{\mathcal{S}}^{n}$ and $Y(x,\theta)\in\mathbb{R}^{m\times n}$, constants $\varepsilon_{i}>0$ ($i=1,2,3$), and SOS polynomials $\lambda_{i}^{l}$ ($i\in\mathbb{Z}_{1}^{{N_{\theta}}},l\in\mathbb{Z}_{1}^{4}$), $\alpha_{i}^{j}~{}(i\in\mathbb{Z}_{1}^{{N_{\theta}}},j\in\mathbb{Z}_{1}^{n})$, $\mu_{j}^{p}$ ($j\in\mathbb{Z}_{1}^{n},p\in\mathbb{Z}_{1}^{3}$) and $\xi_{k}$ ($k\in\mathbb{Z}_{1}^{{n_{\theta}}}$) such that
@@
Part (a): According to Proposition 3, 21a and 21b ensures
@@
Under the dynamics 18 and the control law 23, and the definition of $V$ in 24, we have
@@
where we have leveraged the fact that $\dot{x}=\left({A+BY{X^{-1}}}\right)x$ and $\frac{{\partial{X^{-1}}}}{{\partial{x_{j}}}}=-{X^{-1}}\frac{{\partial X}}{{\partial{x_{j}}}}{X^{-1}}$ [22, Lemma 5] in deriving 27, and $F_{1}(x,\theta,\dot{\theta})$ is defined in 22. According to Proposition 3, 21c implies $F_{1}(x,\theta,\dot{\theta})+\varepsilon_{3}I\leq 0$, $\forall x\in\mathcal{X}$, $\theta\in\mathcal{F}_{\Theta}^{\mathcal{V}}$ with $\mathcal{F}_{\Theta}^{\mathcal{V}}$ defined in 3. Further considering 28, we have
@@
By applying Lyapunov stability theory [21, Theorem 4.10] and considering 26 and 29, we can conclude that under the controller 23, the closed-loop system is locally exponentially stable at the origin, as certified by the Lyapunov function defined in 24.
Part (b): According to Proposition 3, 21d implies ${\footnotesize\begin{bmatrix}1&{C_{i}}X\\ *&X\end{bmatrix}}\geq 0$, $\forall\theta\in\Theta$, which, via Schur complement, further indicates $XC_{i}^{T}{C_{i}}X\leq X$, $\forall\theta\in\Theta$. Multiplying the preceding inequality by $x^{T}X^{-1}$ and its transpose from the left and right, respectively, leads to $|{C_{i}}x|^{2}\leq x^{T}X^{-1}x=V(x,\theta),~{}\forall i\in\mathbb{Z}_{1}^{n},\ \forall\theta\in\Theta$. Thus, $\forall(x,\theta)\in\Omega(V,1)$, we have $|{C_{i}}x|^{2}\leq x^{T}X^{-1}x\leq 1$, ${}\forall i\in\mathbb{Z}_{1}^{n}$, which indicates $x\in\mathcal{X}$. Further considering 29, we can conclude that if $(x(0),\theta(0))\in\Omega(V,1)$, $\theta\in\mathcal{F}_{\Theta}^{\mathcal{V}}$, then
@@
In other words, under the condition that $\theta\in\mathcal{F}_{\Theta}^{\mathcal{V}}$, $\Omega(V,1)$ is an invariant set, and the projection of $\Omega(V,1)$ onto the $x$-plane is a subset of $\mathcal{X}$. As a result, the conditions 29 and 26 indicate $\dot{V}\leq-\frac{\varepsilon_{1}\varepsilon_{3}}{\varepsilon_{2}^{2}}V$, which implies $V(t)\leq V(0)\exp^{-\frac{\varepsilon_{1}\varepsilon_{3}}{\varepsilon_{2}^{2}}t}$, $\forall t\geq 0$. Considering 26, we have $\left\lVert x(t)\right\rVert^{2}\leq\varepsilon_{2}V(t)\leq\varepsilon_{2}V(0)\exp^{-\frac{\varepsilon_{1}\varepsilon_{3}}{\varepsilon_{2}^{2}}t}$, $\forall t\geq 0$, which indicates $\lim_{t\rightarrow 0}x(t)=0$. ∎
@@
Theorem 1 presents the conditions to synthesize a controller to ensure the exponential stability of the closed-loop system, which is stronger than the uniformly asymptotic stability adopted in Sections III and IV. Additionally, condition 5 are convex SOS conditions since they are affine in all decision variables.
@@
Suppose the system 18 is an equivalent representation of the NPV system 1, and all the conditions of Theorem 1 hold. Then the PD Lyapunov function defined in 24 is a PD-CLF for 1, and $\Omega(V,1)$ is a PD-ROS.
@@
Condition 31c ensures $X\geq X_{0}$ and therefore $x^{T}X^{-1}x\leq x^{T}X_{0}^{-1}x$ for any $x\in\mathcal{X},\ \theta\in\Theta$. As a result, $\Omega(x^{T}X_{0}^{-1}x,1)\cap(\mathcal{X}\times\Theta)\subseteq\Omega(V,1)\cap(\mathcal{X}\times\Theta)$.
@@
V-C Handing input constraints
We can further consider input constraints and ensure that the control law in 23 respects such constraints by extending the method in [15, 24] to the parameter-dependent case. Specifically, we consider an input constraint set defined by
$\mathcal{U}\triangleq\{u\in\mathbb{R}^{m}:\left\lvert D_{i}u\right\rvert\leq 1,\ i=1,\dots,N_{u}\},$ (33)
where $D_{i}\in\mathbb{R}^{1\times m}$.
Lemma 5.
Consider an NPV system 18 with the parameters $\theta$ satisfying 2, the set $\mathcal{X}$ defined in 20, and the input constraint set defined in 33. Suppose all the conditions in Theorem 1, i.e., 21a, 21b, 21c and 21d hold, and additionally there exist SOS polynomials $\beta_{i}^{q}$ $(i\in\mathbb{Z}_{1}^{N_{\theta}},q\in\mathbb{Z}_{1}^{m})$ and $\gamma_{j}^{q}$ $(j\in\mathbb{Z}_{1}^{n},q\in\mathbb{Z}_{1}^{m})$ such that
$v_{2}^{T}{\footnotesize\begin{bmatrix}1&D_{j}Y\\ *&X(\tilde{x},\theta)\end{bmatrix}}v_{2}-\sum_{j=1}^{n}\gamma_{j}^{q}c_{j}(x)-\!\sum_{i=1}^{N_{\theta}}\beta_{i}^{q}h_{i}(\theta)\in{\Phi_{\textup{sos}}(x,\theta,v_{2})}\ \forall q\in\mathbb{Z}_{1}^{m}.\!$ (34)
holds. Then, under the control law 23,
$\forall(x(0),\theta(0))\in\Omega(V,1),\ \theta\in\mathcal{F}_{\Theta}^{\mathcal{V}}\Rightarrow u(t)\in\mathcal{U},~{}\forall t\geq 0.$ (35)
Proof.
According to Proposition 3, 34 implies ${\footnotesize\begin{bmatrix}[0.8]1&D_{j}Y\\ *&X\end{bmatrix}}\geq 0$, $\forall j\in\mathbb{Z}_{1}^{m}$, $\forall x\in\mathcal{X}$, which, via Schur complement, further indicates $X^{-1}Y^{T}D_{j}^{T}D_{j}YX^{-1}\leq X^{-1}$, $\forall j\in\mathbb{Z}_{1}^{m}$, $\forall x\in\mathcal{X},\ \theta\in\Theta$. Multiplying the preceding inequality by $x$ and $x^{T}$ from the right and left, respectively, and considering the control law 23, we have $|D_{j}u|^{2}\leq xX^{-1}x$, $\forall j\in\mathbb{Z}_{1}^{m}$, which implies $u(x,\theta)\in\mathcal{U}$, $\forall x\in\mathcal{X},\ \theta\in\Theta$. Further considering 30 leads to 35. ∎
To incorporate input constraints in the synthesis procedure, we just need to add constraints 34 to the SOS optimization problem 32.
VI Simulation Results

### v2

@@ Abstract
This paper introduces the concept of parameter-dependent (PD) control Lyapunov functions (CLFs) for gain-scheduled stabilization of nonlinear parameter-varying (NPV) systems. It shows that given a PD-CLF, a min-norm control law can be constructed by solving a robust quadratic program. For polynomial control-affine NPV systems, it provides convex conditions, based on the sum of squares (SOS) programming, to jointly synthesize a PD-CLF and a PD controller while maximizing the PD region of stabilization. Input constraints can be straightforwardly incorporated into the synthesis procedure. Unlike traditional linear parameter-varying (LPV) methods that rely on linearization or over-approximation to get an LPV model, the proposed framework fully captures the nonlinearities of the system dynamics. The theoretical results are validated through numerical simulations, including a 2D rocket landing case study under varying mass and inertia.
@@ I Introduction
Statement of Contributions: This paper introduces the concepts of parameter-dependent CLF (PD-CLF) for stabilizing NPV systems. It shows that given a PD-CLF, a min-norm control law can be constructed by solving a robust quadratic program and is locally Lipschitz. Additionally, it presents a method for jointly synthesizing a PD-CLF and a PD controller while maximizing the PD region of stabilization based on SOS programming. The theoretical results are validated through numerical simulations, including a 2D rocket landing case study under varying mass and inertia.
@@ II Preliminaries
Unlike existing NPV work [16, 17, 18, 19], we do not require $\dot{\theta}$ to be online measurable. This makes our solution more practical as in the real world it is often challenging to obtain $\dot{\theta}$.
@@ III Parameter-Dependent CLFs for Stabilizing NPV Systems
III-A Constructing min-norm control law via robust quadratic programs
@@
$\alpha_{1}(\left\lVert x\right\rVert)\leq V(x(t),\theta(t))\leq\alpha_{2}(\left\lVert x\right\rVert),\quad\forall t\geq 0,\forall x.$ (14)
@@
III-B Robustness of min-norm control law against input gain uncertainty
An advantage of the min-norm control law in 8 is that it is inherently robust against uncertainty in the input gain. Suppose the original system 1 is perturbed in a way such that the control $u$ is multiplied by a positive gain $\lambda$, i.e., the perturbed system takes the form $\dot{x}=f(x)+g(x)\lambda u$. Then, the min-norm control law constructed using the nominal dynamics will uniformly asymptotically stabilize the perturbed system for all $\lambda>1$; that is, the closed-loop system with the min-norm control law has an infinite gain margin. This observation is formalized in the following lemma.
Lemma 3.
Suppose $V(x,\theta)$ is a PD-CLF for the system 1. Then, under the min-norm control law 8 computed using the nominal dynamics of 1, the origin of the perturbed system $\dot{x}=f(x,\theta)+g(x,\theta)\lambda u$ for all $\lambda>0$ is uniformly asymptotically stable for all admissible trajectories of $\theta$.
Proof.
Note that the CLF derivative condition 8b can be rewritten as 10, repeated below:
$A(\zeta)^{T}u\leq c(\zeta),$ (15)
Now consider the condition
$A(\zeta)^{T}\lambda u\leq c(\zeta),$ (16)
If $c(\zeta)\geq 0$, from 11, we know that the solution for 8 is $u^{*}=0$, which obviously also satisfy 16.
If $c(\zeta)<0$, then $A(\zeta)^{T}u\leq c(\zeta)<0$, which indicates that $A(\zeta)^{T}\lambda u\leq A(\zeta)^{T}u\leq c(\zeta)<0$ for all $\lambda\geq 1$. Notice that 16 is an equivalent reformulation of
$L_{f}V\!+\!L_{g}V(\lambda u)\!+\!\frac{\partial V}{\partial\theta}v\!\leq\!-\alpha(\left\lVert x\right\rVert)$ (17)
Additionally, 5a indicates that 14. By applying Lyapunov stability theory for non-autonomous theorem ([21, Theorem 4.8]) to 13 and 14, we conclude that the perturbed system under the control law in 8 is uniformly asymptotically stable. ∎
III-C Local analysis and PD region of stabilization
We now consider local stabilizability ensured by a PD-CLF.
@@ III-C Local analysis and PD region of stabilization
$\forall(x,\theta,\dot{\theta})\in\mathcal{X}\times\Theta\times\mathcal{V},\quad$
$\hskip 8.53581pt\inf_{u}\underbrace{\frac{\partial V}{\partial x}f+\frac{\partial V}{\partial x}gu+\frac{\partial V}{\partial\theta}\dot{\theta}}_{\triangleq\dot{V}(x,\theta,\dot{\theta},u)}\leq-\alpha(\left\lVert x\right\rVert),\vspace{-8mm}$ (18b)
@@
A PD region of stabilization (PD-ROS) for system 1 is the set of initial states $x(0)$ and parameters $\theta(0)$ from which there exists a control law $K(x,\theta)$ such that $u(t)=K(x(t),\ \forall t\geq 0$, and $\lim_{t\rightarrow 0}x(t)=0$ for all admissible trajectories of $\theta$. Formally, a PD-ROS is a set defined by
$\left\{(x(0),\theta(0))\in\mathbb{R}^{n}\times\Theta\!:\ \forall\theta\in\mathcal{F}_{\Theta}^{\mathcal{V}},\exists u(t)=K(x(t),\theta(t))\ \forall t\geq 0,\textup{ s.t.}\lim_{t\rightarrow 0}x(t)=0.\right\}$
@@
A robust region of stabilization (R-ROS) for system 1 is the set of initial states $x(0)$ from which there exists a control law $K(x,\theta)$ such that $u(t)=K(x(t),\theta(t))$ $\forall t\geq 0$, and $\lim_{t\rightarrow 0}x(t)=0$ for all admissible trajectories of $\theta$. Formally, an R-ROS is a set defined by
$\left\{x(0)\in\!\mathbb{R}^{n}\!\!:\ \forall\theta\in\mathcal{F}_{\Theta}^{\mathcal{V}},\ \exists u(t)=K(x(t),\theta(t))\ \forall t\geq 0,\textup{ s.t.}\lim_{t\rightarrow 0}x(t)=0.\right\}$
@@
Compared to R-ROS, PD-ROS allows us to enlarge the set of states from which the system can be stabilized by explicitly considering $\theta(0)$, the initial value of $\theta$. This is demonstrated by Fig. 2 to be presented in Section V.
@@
Condition 18b ensures that if $(x(0),\theta(0))\in\Omega(V,\rho)$, then there exist $u(t)$ such that
@@
IV Convex synthesis of PD-CLFs Using SOS Programming
@@
Note that a polynomial system 1 can always be written in the form of 1 if $f(x,\theta)$ does not contain terms independent of $x$, i.e., monomials of zero degrees in $x$. When $f(x,\theta)$ contains $x$-independent terms, state transformation may be leveraged to obtain a linear-like form, as we will demonstrate in Section V-B.
@@
IV-A Preliminaries about SOS optimization
@@
(S-procedure) [23] Let $p(x)$ be a a polynomial and let $\mathcal{X}$ be a set defined in 23. Suppose there exist SOS polynomials $\lambda_{i}(x)$, $i=0,1,\dots,n$ such that $\left(1+\lambda_{0}(x)\right)p(x)-\sum_{i=1}^{n}\lambda_{i}(x)c_{i}(x)\in{\Phi_{\textup{sos}}(x)}$. Then, $p(x)\geq 0\textup{ for all }x\in\mathcal{X}$.
@@
[22, Proposition 10] Let $F(x)$ be an $N\times n$ symmetric polynomial matrix of degree $2d$ in $x\in\mathbb{R}^{n}$, $\mathcal{X}$ is a semialgebraic set defined in 23, and $v\in\mathbb{R}^{N}$. Suppose there exist SOS polynomials $\lambda_{i}(x,v)$, $i=1,\dots,n$ such that $v^{T}F(x)v-\sum_{i=1}^{n}\lambda_{i}(x,v)c_{i}(x)\in{\Phi_{\textup{sos}}(x,v)}$. Then, $F(x)\geq 0\textup{ for all }x\in\mathcal{X}$.
IV-B Convex Synthesis of a PD-CLF and a PD Controller
@@
Consider an NPV system 21 with the parameters $\theta$ satisfying 2, and the set $\mathcal{X}$ defined in 23. Suppose there exists polynomial matrices $X(\tilde{x},\theta)\in{\mathcal{S}}^{n}$ and $Y(x,\theta)\in\mathbb{R}^{m\times n}$, constants $\varepsilon_{i}>0$ ($i=1,2,3$), and SOS polynomials $\lambda_{i}^{l}$ ($i\in\mathbb{Z}_{1}^{{N_{\theta}}},l\in\mathbb{Z}_{1}^{4}$), $\alpha_{i}^{j}~{}(i\in\mathbb{Z}_{1}^{{N_{\theta}}},j\in\mathbb{Z}_{1}^{n})$, $\mu_{j}^{p}$ ($j\in\mathbb{Z}_{1}^{n},p\in\mathbb{Z}_{1}^{3}$) and $\xi_{k}$ ($k\in\mathbb{Z}_{1}^{{n_{\theta}}}$) such that
@@
Part (a): According to Proposition 3, 24a and 24b ensures
@@
Under the dynamics 21 and the control law 26, and the definition of $V$ in 27, we have
@@
where we have leveraged the fact that $\dot{x}=\left({A+BY{X^{-1}}}\right)x$ and $\frac{{\partial{X^{-1}}}}{{\partial{x_{j}}}}=-{X^{-1}}\frac{{\partial X}}{{\partial{x_{j}}}}{X^{-1}}$ [22, Lemma 5] in deriving 30, and $F_{1}(x,\theta,\dot{\theta})$ is defined in 25. According to Proposition 3, 24c implies $F_{1}(x,\theta,\dot{\theta})+\varepsilon_{3}I\leq 0$, $\forall x\in\mathcal{X}$, $\theta\in\mathcal{F}_{\Theta}^{\mathcal{V}}$ with $\mathcal{F}_{\Theta}^{\mathcal{V}}$ defined in 3. Further considering 31, we have
@@
By applying Lyapunov stability theory [21, Theorem 4.10] and considering 29 and 32, we can conclude that under the controller 26, the closed-loop system is locally exponentially stable at the origin, as certified by the Lyapunov function defined in 27.
Part (b): According to Proposition 3, 24d implies ${\footnotesize\begin{bmatrix}1&{C_{i}}X\\ *&X\end{bmatrix}}\geq 0$, $\forall\theta\in\Theta$, which, via Schur complement, further indicates $XC_{i}^{T}{C_{i}}X\leq X$, $\forall\theta\in\Theta$. Multiplying the preceding inequality by $x^{T}X^{-1}$ and its transpose from the left and right, respectively, leads to $|{C_{i}}x|^{2}\leq x^{T}X^{-1}x=V(x,\theta),~{}\forall i\in\mathbb{Z}_{1}^{n},\ \forall\theta\in\Theta$. Thus, $\forall(x,\theta)\in\Omega(V,1)$, we have $|{C_{i}}x|^{2}\leq x^{T}X^{-1}x\leq 1$, ${}\forall i\in\mathbb{Z}_{1}^{n}$, which indicates $x\in\mathcal{X}$. Further considering 32, we can conclude that if $(x(0),\theta(0))\in\Omega(V,1)$, $\theta\in\mathcal{F}_{\Theta}^{\mathcal{V}}$, then
@@
In other words, under the condition that $\theta\in\mathcal{F}_{\Theta}^{\mathcal{V}}$, $\Omega(V,1)$ is an invariant set, and the projection of $\Omega(V,1)$ onto the $x$-plane is a subset of $\mathcal{X}$. As a result, the conditions 32 and 29 indicate $\dot{V}\leq-\frac{\varepsilon_{1}\varepsilon_{3}}{\varepsilon_{2}^{2}}V$, which implies $V(t)\leq V(0)\exp^{-\frac{\varepsilon_{1}\varepsilon_{3}}{\varepsilon_{2}^{2}}t}$, $\forall t\geq 0$. Considering 29, we have $\left\lVert x(t)\right\rVert^{2}\leq\varepsilon_{2}V(t)\leq\varepsilon_{2}V(0)\exp^{-\frac{\varepsilon_{1}\varepsilon_{3}}{\varepsilon_{2}^{2}}t}$, $\forall t\geq 0$, which indicates $\lim_{t\rightarrow 0}x(t)=0$. ∎
@@
Theorem 1 presents the conditions to synthesize a controller to ensure the exponential stability of the closed-loop system, which is stronger than the uniformly asymptotic stability adopted in Sections III and III-C. Additionally, condition 5 are convex SOS conditions since they are affine in all decision variables.
@@
Suppose the system 21 is an equivalent representation of the NPV system 1, and all the conditions of Theorem 1 hold. Then the PD Lyapunov function defined in 27 is a PD-CLF for 1, and $\Omega(V,1)$ is a PD-ROS.
@@
Condition 34c ensures $X\geq X_{0}$ and therefore $x^{T}X^{-1}x\leq x^{T}X_{0}^{-1}x$ for any $x\in\mathcal{X},\ \theta\in\Theta$. As a result, $\Omega(x^{T}X_{0}^{-1}x,1)\cap(\mathcal{X}\times\Theta)\subseteq\Omega(V,1)\cap(\mathcal{X}\times\Theta)$.
@@
IV-C QP-based control law vs pre-defined control law
Since the resulting Lyapunov function from solving 35 is a PD-CLF according to Lemma 5, we can again construct the control signal by solving a QP similar to 8. From 32, the QP becomes
$u^{\ast}(x,\theta)=\min_{u}\ u^{T}H(\theta)u$ (36a)
s.t. $L_{f}V\!+\!L_{g}Vu\!+\!\frac{\partial V}{\partial\theta}v\leq-\varepsilon_{3}x^{T}X^{-1}X^{-1}x.$ (36b)
where we also introduce a $\theta$-dependent positive definite matrix $H(\theta)$ to make the cost function more general. Notice that can be replaced by $L_{f}V\!+\!L_{g}Vu\!+\!\frac{\partial V}{\partial\theta}v\leq-\frac{\varepsilon_{3}}{\varepsilon_{2}^{2}}\left\lVert x\right\rVert^{2},\ \forall v\in\textup{ver}(\mathcal{V})$, which is a more relaxed condition and will lead to more conservative behavior.
Remark 8.
The QP problem 36 is guaranteed to be feasible for any $(x,\theta)\in\Omega(V,1)$ due to Theorem 1.
Since an input control law defined in 26 is simultaneously obtained when solving 35 to obtain the PD-CLF, one may wonder why we still need to resort to the QP 36 to construct the min-norm control law. The min-norm may be preferred when avoiding input saturation is a concern, as min-norm control law typically generates small control signal as demonstrated in case studies in Section V). Additionally, min-norm control law may improve the robustness of the closed-loop system against input uncertainty, as demonstrated in [25].
V Simulation Results